what is the kinetic energy of a 3-kg object moving at 7 m/s ?
1
Expert's answer
2013-01-15T08:41:24-0500
T = m*v^2/2
T - kinetic energy of the body
m - mass of the body
v - speed of the body
In our case:
v = 7 m/s
m = 3 kg
Therefore
T = 3kg*(7m/s)^2/2 = 147/2 J = 73.5 J